# PointAD+: Learning Hierarchical Representations for Zero-shot 3D Anomaly Detection

### v2

@@ 4.3 Result Analysis
PointAD+ characterizes itself by capturing abnormality from two aspects: implicit 3D abnormality where point anomalies are projected into the pixel space; (2) explicit 3D abnormality, which reveals the inherent point relations. In this section, we provide a more in-depth analysis of how they contribute to the success of PointAD+.
@@ 4.3.1 Implicit point learning on rendering layer
Figure 7: Visualization of explicit and implicit point score maps.
@@ 4.3.2 Explicit point learning on geometric layer
We further study what PointAD+ learns from the geometry layer. The top row of Figure 7 presents explicit 3D segmentation according to Eq. 5. It can be observed that explicit point learning effectively captures point-level abnormality in unseen objects. By incorporating geometric information, explicit point learning has a stronger ability to detect abnormal point relations, making it more accurate in identifying such anomalies compared to implicit point learning. As shown in the four images on the left, explicit point segmentation successfully localizes anomaly regions with high precision. However, implicit point learning, which directly aggregates pixel abnormality into point abnormality, sometimes encounters challenges in such cases, as demonstrated in the implicit point segmentation described in Eq. 3. This does not necessarily mean that explicit point learning outperforms implicit point learning in all scenarios. The four right images depict that explicit point learning encounters challenges with edge points, where rapid changes in point relations can lead to false anomaly alarms. In contrast, implicit point segmentation, by solely aggregating rendering pixel anomalies, naturally circumvents this issue. Therefore, PointAD+ proposes hierarchical representation learning to complement explicit and implicit point learning, enabling a more holistic understanding of abnormality.
(a) Visualization comparison on ZS M3D between (a) with hybrid loss and (b) without 2D glocal loss.
@@
PointAD+ and PointAD are unified frameworks that can not only capture point anomalies but also handle 2D information in a plug-and-play manner. To study how it works, we visualize the M3D results of PointAD on MVTec3D-AD in Figure LABEL:fig6:_hybrid_loss. The surface damage on the potato presents a similar appearance to the object foreground, which makes it difficult to detect this anomaly with RGB information. On the contrary, the point relations for the color stain on foam are the same as those of normal, but they have a clear distinction in the RGB information. PointAD can integrate these two modalities, thereby complementing their respective advantages. We further investigate the reason why PointAD+ and PointAD can directly leverage both modalities. For this purpose, we experiment without 2D glocal loss. As shown in Figure LABEL:fig6:_hybrid_loss_without_2D_glocal_loss, without 2D glocal loss, significant noise disrupts and even covers the RGB score maps, resulting in unpromising multimodal fusions. This illustrates the importance of explicit constraints on the 2D space. Hence, we conclude that the robust multimodal detection capability of our model stems from the collaboration optimization in both 3D and 2D spaces during training.
@@ 4.3.4 Complexity analysis
TABLE V: Comparison of computation overhead with SOTA approaches on MVTec3D-AD.
Methods
Inference
time (s)
FPS
GPU memory
usage (Peak)
Point detection
Local Global
BTF (Un.) 0.18 5.56 1934MB (97.6, 92.3) (76.3, 91.8)
SDM (Un.) 0.14 7.14 2716MB (97.0, 91.8) (96.7, 90.9)
M3DM (Un.) 2.86 0.35 7494MB (92.8, 91.6) (85.6, 93.4)
PointCLIP V2 (ZS) 1.52 0.66 9747MB (87.4, 52.3) (78.3, 49.4)
CLIP + Rendering (ZS) 0.27 3.61 3685MB (61.2, 85.8) (-, 54.4)
Cheraghian (ZS) 0.35 2.86 4847MB (88.2, 57.0) (53.6, 81.7)
AnomalyCLIP (ZS) 0.19 5.26 3348MB (88.9, 60.9) (56.4, 83.5)
PointAD (ZS) 0.40 2.52 4275MB (95.5, 84.4) (82.0, 94.2)
PointAD+ (ZS) 0.46 2.18 5811MB (96.9, 88.7) (82.5, 94.8)
TABLE VI: Ablation on the proposed modules. F stage and S stage. represent the first and second stages of G-aggregation, respectively.
$L^{r}$ $L^{g}$ $L^{cross}$ Point detection Multimodal detection
$L_{I-3d}^{global}$ $L_{I-3d}^{local}$ $L_{I-2d}^{global}$ $L_{I-2d}^{local}$ F stage S stage Local Global Local Global
@@ Methods
TABLE VII: Ablation on the view number.
View number Point detection Multimodal detection
@@
11 (96.6, 87.4) (82.7, 94.4) (97.9, 92.3) (87.5, 96.5)
TABLE VIII: Ablation on learnable prompt length.
Length of learnable prompt Point detection Multimodal detection
@@
6 (96.0, 88.2) (82.4, 94.1) (97.9, 92.4) (87.2, 96.3)
8 (96.2, 88.1) (82.2, 94.4) (98.1, 92.3) (87.5, 96.3)
10 (96.5, 88.5) (82.7, 94.5) (98.1, 92.6) (87.5, 96.3)
12 (96.9, 88.7) (82.5, 94.8) (98.2, 93.1) (87.7, 96.6)
14 (96.7, 88.3) (82.3, 94.4) (98.0, 92.8) (87.3, 96.5)
TABLE IX: Density ablation.
Downsample Point detection
ratio Local Global
100% (96.9, 88.7) (82.5, 94.8)
70% (96.2, (86.1) (82.8, 94.5)
50% (95.8, 85.1) (81.2, 93.8)
30% (95.7, 84.3) (79.0, 92.8)
20% (95.5, 83.5) (75.0, 91.5)
TABLE X: Lighting ablation.
Lighting Point detection
Local Global
++ (96.9, 88.3) (82.1, 94.8)
+ (96.9, 88.3) (82.0, 95.1)
original (96.9, 88.7) (82.5, 94.8)
- (96.7, 87.9) (81.8, 94.8)
(96.7, 87.7) (81.5, 94.6)
The computation overhead is also an indicator to assess the effectiveness of the proposed method. We compare the computation overhead in Table V, including inference time per image, frames per second (FPS), and GPU memory consumption with a batch size of 1. For a fair comparison, we maintain NVIDIA RTX 3090 24GB GPU idle. Un. denotes the unsupervised setting, and ZS denotes the zero-shot setting. The performance and computational overhead of PointAD and PointAD+ are comparable to those of unsupervised methods. As for zero-shot detection, compared to PointCLIP V2, PointAD requires less time to infer an image, achieving a higher 2.52 vs. 0.66 FPS with lower graphic memory usage. PointAD+ needs a slightly higher computation overhead over PointAD, i.e., 2.52 vs. 2.18 FPS, and attains a noticeable performance improvement on ZS 3D and M3D. Vanilla CLIP + rendering has the fastest inference speed, but its detection performance lags behind PointAD+ and PointAD. Therefore, PointAD+ achieves a favorable trade-off between performance and overhead.
@@ 5 Ablation Study
Module ablation. We evaluate the effectiveness of the proposed techniques by progressively incorporating the introduced modules. As shown in Table VI, the baseline vanilla model, which represents CLIP + rendering, performs poorly on both 3D and M3D anomaly detection due to CLIP’s focus on aligning 2D object semantics rather than anomaly semantics. Adding the 3D global branch incorporates global anomaly semantics into PointAD+, improving both local and global anomaly detection performance. Although the 3D local branch further enhances performance, local accuracy on M3D detection declines due to the absence of 2D constraints, leading to inefficient multimodal fusion when integrating 2D RGB information in a plug-and-play manner. The inclusion of the 2D global branch explicitly introduces 2D anomaly information, enabling PointAD+ to achieve overall performance gains and better multimodal fusion. By focusing on 2D anomaly regions, PointAD+ develops a deeper understanding of point clouds from 2D representations, improving multimodal segmentation performance from 92.9% to 97.2% P-AUROC and from 84.4% to 90.2% AUPRO. So far, PointAD+ does not explicitly account for point relationships within point clouds. To address this, explicit point learning is introduced to model abnormal spatial relationships. PointAD+ leverages geometric characteristics, achieving notable improvements in segmentation performance. Specifically, P-AUROC increases from 95.5% to 96.3% and AUPRO rises from 84.4% to 87.5% on ZS 3D detection. On M3D detection, P-AUROC improves from 97.2% to 97.9% and AUPRO increases from 90.2% to 92.5%. Finally, cross-alignment is introduced to enhance interactions between hierarchical representations, leading to further performance gains over these metrics.
@@
Ablation on view number. PointAD+ interprets point clouds through 2D renderings, where the number of rendering views directly influences the extent of 3D information captured. As shown in Table VII, an excessive number of views introduces redundancy and noise, hindering accurate point representation. Conversely, an insufficient number of views results in information loss. Therefore, selecting an optimal number of views enhances point cloud understanding by effectively leveraging informative perspectives.
@@
Ablation on length of learnable prompts. Here, we examine the impact of the length of learnable text prompt templates. As shown in Table VIII, increasing the length of word embeddings enhances PointAD+’s capacity to incorporate both 3D and 2D anomaly semantics, contributing to the performance gain. However, further increasing the length (e.g., from 12 to 14) does not yield additional performance gains and instead results in a declining trend. Excessively long prompts may introduce noise, which disrupts the detection process. This demonstrates that both excessive and insufficient lengths of learnable word embeddings can degrade performance. In our study, an optimal prompt length of 12 enables PointAD+ to achieve comprehensive performance in both 3D and M3D anomaly detection.
@@
Ablation on input density. Further exploration of the robustness of PointAD+ under varying conditions could enhance its generalized detection performance. We investigate the impact of input point density on the performance of PointAD+ and PointAD. To create low-resolution point clouds, we downsample the original high-resolution point clouds using Farthest Point Sampling (FPS) with various sampling ratios. This generates corresponding low-resolution datasets for evaluating PointAD+. Visualizations of these datasets are provided in Figure LABEL:fig:_Visualization_with_different_resolutions. We train PointAD+ on the original dataset and evaluate its performance on point clouds with varied resolutions. As shown in Table IX, PointAD+ demonstrates robust detection capabilities across a wide range of resolutions, including downsampling ratios of 20%, 30%, 50%, and 70%. Notably, PointAD+ achieves SOTA performance even at 20% resolution, indicating that it is highly adaptable to point clouds with varying resolutions.
@@
Ablation on rendering lighting. In addition to evaluating the input resolution of renderings, we assess the robustness of PointAD+ to varying lighting conditions. To examine the impact of rendering lighting, we adjusted the lighting settings used to render point clouds in Open3D, generating a series of datasets with a range of lighting intensities. Specifically, we applied both stronger and weaker lighting compared to the original dataset, covering a broad range of lighting conditions. We denote stronger and the strongest lighting levels as "+" and "++," respectively, and weaker and the weakest lighting levels as "-" and "–". Figure LABEL:fig:_Visualization_with_different_rendering_lighting. presents the visualizations of the resulting samples on MVTec3D-AD, where PointAD+ was trained on the original dataset and tested on these lighting-variant datasets. As shown in Table X, PointAD+ consistently detects anomalies despite significant variations in rendering lighting, demonstrating its robustness to such discrepancies.

### v3

@@ 4.3 Result Analysis
PointAD+ characterizes itself by capturing abnormality from two aspects: implicit 3D abnormality where point anomalies are projected into the pixel space; explicit 3D abnormality, which reveals the inherent point relations. In this section, we provide a more in-depth analysis of how they contribute to the success of PointAD+.
@@ 4.3.2 Explicit point learning on geometric layer
We further examine what PointAD+ learns from the geometry layer according to Eq. 15. As shown in Figure LABEL:fig:_ei1, explicit point learning effectively captures point-level abnormalities in previously unseen objects. By incorporating geometric information, it demonstrates a stronger ability to detect abnormal point relations, making it more accurate than implicit point learning in identifying such anomalies. However, in Figure LABEL:fig:_ei2, when point relations change rapidly, explicit point representations may produce false anomaly alarms.
(a) Visualization comparison on ZS multimodal 3D between (a) with hybrid loss and (b) without 2D glocal loss.
@@
PointAD+ and PointAD are unified frameworks that can not only capture point anomalies but also handle 2D information in a plug-and-play manner. To study how it works, we visualize the multimodal 3D results of PointAD on MVTec3D-AD in Figure LABEL:fig6:_hybrid_loss. The surface damage on the potato presents a similar appearance to the object foreground, which makes it difficult to detect this anomaly with RGB information. On the contrary, the point relations for the color stain on foam are the same as those of normal, but they have a clear distinction in the RGB information. PointAD can integrate these two modalities, thereby complementing their respective advantages. We further investigate the reason why PointAD+ and PointAD can directly leverage both modalities. For this purpose, we experiment without 2D glocal loss. As shown in Figure LABEL:fig6:_hybrid_loss_without_2D_glocal_loss, without 2D glocal loss, significant noise disrupts and even covers the RGB score maps, resulting in unpromising multimodal fusions. This illustrates the importance of explicit constraints on the 2D space. Hence, we conclude that the robust multimodal detection capability of our model stems from the collaboration optimization in both 3D and 2D spaces during training.
TABLE V: Analysis on aggregation strategy $\ell$.
$\ell$ Point detection Multimodal detection
Local Global Local Global
Mean (96.9, 88.7) (82.5, 94.8) (98.2, 93.1) (87.7, 96.6)
Max (97.0, 88.3) (82.5, 94.4) (97.9, 93.5) (87.8, 96.4)
Parameteric network (97.1, 88.7) (82.4, 94.4) (98.0, 93.2) (88.0, 96.5)
TABLE VI: Comparison of computation overhead with SOTA approaches on MVTec3D-AD.
Methods GFLOPs FPS GPU memory usage (Peak) Point detection
Local Global
PointCLIP V2 (ZS) 1063.80 0.66 9747MB (87.4, 52.3) (78.3, 49.4)
CLIP + Rendering (ZS) 1057.36 3.61 3685MB (61.2, 85.8) (-, 54.4)
Cheraghian (ZS) 31.49 2.86 4847MB (88.2, 57.0) (53.6, 81.7)
AnomalyCLIP (ZS) 1656.34 5.26 3348MB (88.9, 60.9) (56.4, 83.5)
PointAD (ZS) 1656.34 2.52 4275MB (95.5, 84.4) (82.0, 94.2)
PointAD+ (ZS) 1658.34 2.18 5811MB (96.9, 88.7) (82.5, 94.8)
4.3.4 Aggregation strategy
By default, we adopt the mean aggregation strategy to obtain explicit point representations. We also investigate two alternative strategies: max strategy, which selects the maximum pixel representation across different views as the corresponding explicit point representation; network-parametric strategy, which employs an MLP followed by a softmax layer to compute aggregation weights of the pixel representations. As shown in Table V, PointAD+ is not restricted to a specific aggregation scheme but generalizes across different strategies. We adopt the mean strategy as it provides a favorable trade-off between performance and computational overhead.
@@ 4.3.5 Complexity analysis
The computation overhead is also an indicator to assess the effectiveness of the proposed method. We compare the computation overhead in Table VI, including GFLOPs, frames per second (FPS), and GPU memory consumption with a batch size of 1. For a fair comparison, we maintain NVIDIA RTX 3090 24GB GPU idle. PointAD+ requires only about 2.0 more GFLOPs than PointAD (1658.34 vs. 1656.34) yet achieves a significant performance improvement. In contrast, other CLIP-based methods such as AnomalyCLIP and PointCLIP V2 consume comparable or even fewer GFLOPs but deliver worse results. Although CLIP+Rendering and Cheraghian use fewer GFLOPs, they fail to provide competitive performance. The performance and computational overhead of PointAD and PointAD+ are comparable to those of unsupervised methods. As for ZS detection, compared to PointCLIP V2, PointAD requires less time to infer an image, achieving a higher 2.52 vs. 0.66 FPS with lower graphic memory usage. PointAD+ needs a slightly higher computation overhead over PointAD, i.e., 2.52 vs. 2.18 FPS, and attains a noticeable performance improvement on ZS 3D and multimodal 3D. CLIP + rendering has the fastest inference speed, but its detection performance lags behind PointAD+ and PointAD. Therefore, PointAD+ achieves a favorable trade-off between performance and overhead.
TABLE VII: Ablation on proposed modules. F stage and S stage represent the first and second stages of G-aggregation, respectively.
$L^{r}$ $L^{g}$ $L^{c}$ Point detection Multimodal detection
$L_{I(3d)}^{g}$ $L_{I(3d)}^{l}$ $L_{I(2d)}^{g}$ $L_{I(2d)}^{l}$ F S Local Global Local Global
@@
TABLE VIII: Ablation on the neighbors $k$.
k Point detection Multimodal detection
@@
15 (97.0, 88.3) (82.6, 94.4) (97.9, 93.2) (87.8, 96.4)
20 (96.9, 88.5) (82.5, 94.5) (97.8, 93.5) (87.6, 96.5)
25 (96.9, 88.4) (82.5, 94.5) (97.8, 93.1) (87.6, 96.4)
TABLE IX: Ablation on the view number. The absolute numbers report PointAD+ performance. Arrows denote relative changes against baseline PointAD: for improvements, purple for degradations.
K Point detection Multimodal detection
@@
[…]
(a) The effect of $\alpha$ to the overall performance.
@@ 5 Ablation Study
TABLE X: Prompt length ablation.
E Point detection Multimodal detection
Local Global Local Global
[…]
TABLE XI: Density ablation.
Down sample Point detection Multimodal detection
Local Global Local Global
[…]
TABLE XII: Lighting ablation.
Lighting Point detection Multimodal detection
Local Global Local Global
[…]
TABLE XIII: Occlusion ablation.
Method Point detection Multimodal detection
Local Global Local Global
[…]
TABLE XIV: Quality ablation.
Blur sample Point detection Multimodal detection
Local Global Local Global
[…]
TABLE XV: Angle ablation.
Angle shift Point detection Multimodal detection
Local Global Local Global
[…]
5.1 Ablation on modules and key hyperparameters
@@
Module ablation. We evaluate the effectiveness of the proposed techniques by progressively incorporating the introduced modules. As shown in Table VII, the baseline vanilla model, which represents CLIP + rendering, performs poorly on both 3D and multimodal 3D anomaly detection due to CLIP’s focus on aligning 2D object semantics rather than anomaly semantics. Adding the 3D global branch incorporates global anomaly semantics into PointAD+, improving both local and global anomaly detection performance. Although the 3D local branch further enhances performance, local accuracy on multimodal 3D detection declines due to the absence of 2D constraints, leading to inefficient multimodal fusion when integrating 2D RGB information in a plug-and-play manner. The inclusion of the 2D global branch explicitly introduces 2D anomaly information, enabling PointAD+ to achieve overall performance gains and better multimodal fusion. By focusing on 2D anomaly regions, PointAD+ develops a deeper understanding of point clouds from 2D representations, improving multimodal segmentation performance from 92.9% to 97.2% P-AUROC and from 84.4% to 90.2% AUPRO. So far, PointAD+ does not explicitly account for point relationships within point clouds. To address this, explicit point learning is introduced to model abnormal spatial relationships. PointAD+ leverages geometric characteristics, achieving notable improvements in segmentation performance. Specifically, P-AUROC increases from 95.5% to 96.3% and AUPRO rises from 84.4% to 87.5% on ZS 3D detection. On multimodal 3D detection, P-AUROC improves from 97.2% to 97.9% and AUPRO increases from 90.2% to 92.5%. Finally, cross-alignment is introduced to enhance interactions between hierarchical representations, leading to further performance gains over these metrics.
Ablation on $\alpha$. The coefficient $\alpha$ balances the feature fusion between the F and S stages. We investigate the effect of the coefficient $\alpha$ on the model’s performance. We set $\alpha\in\{0,0.1,0.3,0.5,0.7,0.9,1\}$. As shown in Figure 10(a), increasing $\alpha$ initially improves both image-/pixel- auroc performance. However, as $\alpha$ continues to increase, performance begins to decline. Empirically, $\alpha=0.5$ achieves a good balance between image-level and pixel-level performance.
Ablation on the number of neighbors The number of neighbors $k$ determines the receptive field of each point within its local neighborhood. As illustrated in Table VIII, we further explore $k\in\{5,10,15,20,25\}$ to study its impact. For point detection, the best local accuracy is achieved at $k=10$ (96.9 AUROC and 88.7 AUPRO), and the best global accuracy is also observed at $k=10$ (82.5 AUROC and 94.8 AP). For multimodal detection, both $k=10$ and $k=20$ yield superior performance. Overall, the detection results remain stable across different values of $k$. This indicates that PointAD+ is robust to the number of neighboring points..
@@
Ablation on view number. PointAD+ interprets point clouds through 2D renderings, where the number of rendering views directly influences the extent of 3D information captured. As shown in Table IX, an excessive number of views introduces redundancy and noise, hindering accurate point representation. Conversely, an insufficient number of views results in information loss. Therefore, selecting an optimal number of views enhances point cloud understanding by effectively leveraging informative perspectives.
@@
Ablation on length of learnable prompts. Here, we examine the impact of the length of learnable text prompt templates. As shown in Table X, increasing the length of word embeddings enhances PointAD+’s capacity to incorporate both 3D and 2D anomaly semantics, contributing to the performance gain. However, further increasing the length (e.g., from 12 to 14) does not yield additional performance gains and instead results in a declining trend. Excessively long prompts may introduce noise, which disrupts the detection process. This demonstrates that both excessive and insufficient lengths of learnable word embeddings can degrade performance. In our study, an optimal prompt length of 12 enables PointAD+ to achieve comprehensive performance in both 3D and multimodal 3D anomaly detection.
5.2 Ablation on the varied rendering conditions.
PointAD+ interprets point clouds based on the rendering representations through CLIP. Here, we investigate the effect of the rendering condition, including point density in Table XI, lighting ablation in Table XII, point occlusion in Table XIII, rendering quality in Table XIV, and rendering angle in Table XV.
@@
Ablation on input density. We investigate the impact of input point density on the performance of PointAD+ and PointAD. To create low-resolution point clouds, we downsample the original high-resolution point clouds using Farthest Point Sampling (FPS) with various sampling ratios. This generates corresponding low-resolution datasets for evaluating PointAD+. Visualizations of these datasets are provided in Figure LABEL:fig:_Visualization_with_different_resolutions. We train PointAD+ on the original dataset and evaluate its performance on point clouds with varied resolutions. As shown in Table XI, PointAD+ demonstrates robust detection capabilities across a wide range of resolutions, including downsampling ratios of 20%, 30%, 50%, and 70%. Notably, PointAD+ achieves SOTA performance even at 20% resolution, indicating that it is highly adaptable to point clouds with varying resolutions.
@@
Ablation on rendering lighting. To examine the impact of rendering lighting, we adjusted the lighting settings used to render point clouds in Open3D, generating a series of datasets with a range of lighting intensities. Specifically, we applied both stronger and weaker lighting compared to the original dataset, covering a broad range of lighting conditions. We denote stronger and the strongest lighting levels as "+" and "++," respectively, and weaker and the weakest lighting levels as "-" and "–". Figure LABEL:fig:_Visualization_with_different_rendering_lighting. presents the visualizations of the resulting samples on MVTec3D-AD, where PointAD+ was trained on the original dataset and tested on these lighting-variant datasets. As shown in Table XII, PointAD+ consistently detects anomalies despite significant variations in rendering lighting, demonstrating its robustness to such discrepancies.
Ablation on rendering angles. To further examine the effect of the absolute angle, we adjust the rendering orientations while preserving the relative angular discrepancy. In our setup, the baseline adjacent angle discrepancy is $\tfrac{1}{5}\pi$. This gap is subdivided into four intervals, and we apply angular shifts of $\tfrac{1}{20}\pi$, $\tfrac{2}{20}\pi$, and $\tfrac{3}{20}\pi$ to evaluate the influence of different orientation changes. As reported in Table XV, PointAD+ delivers stable performance under these variations and maintains the superiority of PointAD. This confirms its robustness to unseen rendering angles.
Ablation on rendering quality. In our experiments, we relied on the Open3D Library to generate point cloud renderings, though it does not provide an API for explicitly controlling rendering quality. we evaluated PointAD on MVTec3D-AD under blur levels $\sigma\in\{1,5,9\}$. As summarized in Table XIV, the detection performance of PointAD+ decreases as the rendering quality degrades (i.e., with larger $\sigma$ values). Nevertheless, even under heavy blur ($\sigma=9$), the model maintains acceptable accuracy and consistently outperforms baselines that rely on high-quality renderings.
Ablation on point occlusion. We next evaluated detection performance under occluded point clouds by removing points invisible from a given rendering angle and projecting the remaining ones. In some cases, abnormal regions were completely hidden, effectively converting anomalies into normal instances. To reduce this effect, we selected angles that preserved visibility of anomalous regions. Table XIII shows that PointAD+ exhibits performance degradation when tested on the occluded point clouds. We attribute this degradation to two factors: first, occluded point clouds may lose part of the anomaly semantics, making anomalies harder to recognize; second, occlusion can create artificial sinkholes on the surface, which may cause PointAD+ to mistakenly identify these regions as hole anomalies. Although PointAD+ outperforms PointAD, its performance gain on the occluded dataset is smaller than on the original dataset. We analyze that PointAD+ leverages intrinsic point relations. When points in certain regions are completely removed, these relations are disrupted and hinder geometric modeling.